OpenStudy (anonymous):

What is the range of this function? f(x) = log(1/4) x I know the domain is all real numbers..

OpenStudy (anonymous):

\[\log_{\frac{1}{4}}x\]??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Hmm Range should be \[ y = \log_{\frac{1}{4}}x\] \[\left(\frac{1}{4}\right)^{y} = x\] Y is valid for all real numbers

OpenStudy (anonymous):

And Yeah Domain isn't all real numbers, domain is x>0

OpenStudy (anonymous):

oh..

OpenStudy (anonymous):

Yup :-)

OpenStudy (anonymous):

so the range is all real numbers?

OpenStudy (anonymous):

Yeah!

OpenStudy (anonymous):

ok thanks